# AICC: Parse HTML Finer, Make Models Better —— A 7.3T AI-Ready Corpus Built by a Model-Based HTML Parser

### v1

@@ Abstract
We introduce MinerU-HTML, a novel extraction pipeline that reformulates content extraction as a sequence labeling problem solved by a 0.6B-parameter language model. Unlike text-density heuristics, MinerU-HTML leverages semantic understanding and employs a two-stage formatting pipeline that explicitly categorizes semantic elements before converting to Markdown. Crucially, its model-based approach is inherently scalable, whereas heuristic methods offer limited improvement pathways. On MainWebBench, our benchmark of 7,887 annotated web pages, MinerU-HTML achieves 81.82% ROUGE-N F1 compared to Trafilatura’s 63.58%, with exceptional structured element preservation (90.93% for code blocks, 93.99% for formulas).
Using MinerU-HTML, we construct AICC (AI-ready Common Crawl), a 7.3-trillion token multilingual corpus from two Common Crawl snapshots. In controlled pretraining experiments where AICC and Trafilatura-extracted TfCC undergo identical filtering, models trained on AICC (62B tokens) achieve 50.82% average accuracy across 13 benchmarks, outperforming TfCC by 1.08pp—providing direct evidence that extraction quality significantly impacts model capabilities. AICC also surpasses RefinedWeb and FineWeb on key benchmarks. We publicly release MainWebBench, MinerU-HTML, and AICC, demonstrating that HTML extraction is a critical, often underestimated component of web corpus construction.
@@ 1 Introduction
    Extraction Quality Matters: On MainWebBench, our newly constructed benchmark of 7,887 annotated web pages, MinerU-HTML achieves a ROUGE-N F1 score of 0.8182, substantially outperforming Trafilatura (0.6358). For structured element preservation, MinerU-HTML achieves 0.9093 edit similarity for code blocks and 0.9399 for formulas, compared to 0.1305 and 0.6107 for Trafilatura.
@@
Our work makes the following contributions: (1) We introduce MinerU-HTML, a semantic-aware HTML extraction pipeline that significantly outperforms existing heuristic methods on both content fidelity and structured element preservation. Critically, MinerU-HTML’s model-based approach offers inherent scalability advantages over rule-based extractors, as its performance can improve with more data and stronger base models (Figure 2). (2) We construct and publicly release MainWebBench, a comprehensive benchmark for evaluating main content extraction and structured element preservation with 7,887 annotated pages. (3) We release AICC, a large-scale pretraining corpus demonstrating that high-quality extraction can rival or exceed the benefits of aggressive filtering strategies for downstream model performance. (4) We provide extensive pretraining experiments isolating the contribution of extraction quality, shifting attention to this critical but often overlooked component of web data curation.
@@ 3 Evaluation of MinerU-HTML
To comprehensively evaluate MinerU-HTML, we construct MainWebBench, a new benchmark designed to assess both main content extraction and structured element preservation. We compare MinerU-HTML against established baseline methods on MainWebBench and validate its generalization capabilities on the external WCEB benchmark.
3.1 MainWebBench
@@ 3.1.1 Dataset Overview
MainWebBench comprises 7,887 carefully annotated web pages, each containing: (1) the raw HTML document, (2) ground-truth Main-HTML identified by human annotators as a valid DOM subtree, (3) the corresponding Markdown representation, and (4) rich metadata including language, content type, difficulty level, and structural characteristics. This comprehensive annotation enables fine-grained performance analysis across diverse web content types. An example data entry is shown in Appendix Figure 9.
@@ 3.1.3 WebMainBench-Structured: A Subset for Structured Element Extraction
To rigorously evaluate structured element preservation during main content extraction, we curate WebMainBench-Structured, a focused subset of 545 pages selected from MainWebBench that are rich in mathematical formulas, code blocks, and tables—elements critical for technical and educational content but often corrupted by traditional extraction methods. In addition to standard annotations, each page includes a groundtruth_content field containing ground-truth Markdown text with all structured elements faithfully preserved by human annotators, enabling precise element-level evaluation.
@@ 3.3.1 Main Content Extraction
Table 1: Mean ROUGE-N F1 on MainWebBench with different tracks
@@ 3.4 Generalization to WCEB
Table 3 shows that MinerU-HTML maintains strong performance on WCEB, achieving an overall score of 0.8002 and outperforming the strongest baseline, Trafilatura (0.7833). The strong performance across both MainWebBench and WCEB validates MinerU-HTML’s robustness and generalization capabilities across diverse web content.
@@ 5 Conclusion
We release MinerU-HTML, MainWebBench, and AICC to facilitate further research into semantic-aware extraction methods and their impact on pretraining data quality.
@@ 6.2 Document Formatting
Figure 9: An example data from MainWebBench. It includes the raw source, the ground-truth Main-HTML, its Markdown conversion, and a rich set of metadata for fine-grained analysis.
@@ 7.1 Benchmark Construction
Data Sampling. MainWebBench is constructed using a hybrid sampling strategy to ensure both broad representation and relevance. 90% of the samples are randomly drawn from the Common Crawl dataset to cover the long-tail web, while the remaining 10% are sampled from a list of top-ranking websites (Chinaz Alexa [^4] https://malexa.chinaz.com/) to include popular, professionally designed pages. The final benchmark is highly diverse, containing pages from 5,434 unique top-level and 5,904 unique second-level domains.

### v2

@@ Abstract
We introduce MinerU-HTML, a novel extraction pipeline that reformulates content extraction as a sequence labeling problem solved by a 0.6B-parameter language model. Unlike text-density heuristics, MinerU-HTML leverages semantic understanding and employs a two-stage formatting pipeline that explicitly categorizes semantic elements before converting to Markdown. Crucially, its model-based approach is inherently scalable, whereas heuristic methods offer limited improvement pathways. On WebMainBench, our benchmark of 7,887 annotated web pages, MinerU-HTML achieves 81.82% ROUGE-N F1 compared to Trafilatura’s 63.58%, with exceptional structured element preservation (90.93% for code blocks, 93.99% for formulas).
Using MinerU-HTML, we construct AICC (AI-ready Common Crawl), a 7.3-trillion token multilingual corpus from two Common Crawl snapshots. In controlled pretraining experiments where AICC and Trafilatura-extracted TfCC undergo identical filtering, models trained on AICC (62B tokens) achieve 50.82% average accuracy across 13 benchmarks, outperforming TfCC by 1.08pp—providing direct evidence that extraction quality significantly impacts model capabilities. AICC also surpasses RefinedWeb and FineWeb on key benchmarks. We publicly release WebMainBench, MinerU-HTML, and AICC, demonstrating that HTML extraction is a critical, often underestimated component of web corpus construction.
@@ 1 Introduction
    Extraction Quality Matters: On WebMainBench, our newly constructed benchmark of 7,887 annotated web pages, MinerU-HTML achieves a ROUGE-N F1 score of 0.8182, substantially outperforming Trafilatura (0.6358). For structured element preservation, MinerU-HTML achieves 0.9093 edit similarity for code blocks and 0.9399 for formulas, compared to 0.1305 and 0.6107 for Trafilatura.
@@
Our work makes the following contributions: (1) We introduce MinerU-HTML, a semantic-aware HTML extraction pipeline that significantly outperforms existing heuristic methods on both content fidelity and structured element preservation. Critically, MinerU-HTML’s model-based approach offers inherent scalability advantages over rule-based extractors, as its performance can improve with more data and stronger base models (Figure 2). (2) We construct and publicly release WebMainBench, a comprehensive benchmark for evaluating main content extraction and structured element preservation with 7,887 annotated pages. (3) We release AICC, a large-scale pretraining corpus demonstrating that high-quality extraction can rival or exceed the benefits of aggressive filtering strategies for downstream model performance. (4) We provide extensive pretraining experiments isolating the contribution of extraction quality, shifting attention to this critical but often overlooked component of web data curation.
@@ 3 Evaluation of MinerU-HTML
To comprehensively evaluate MinerU-HTML, we construct WebMainBench, a new benchmark designed to assess both main content extraction and structured element preservation. We compare MinerU-HTML against established baseline methods on WebMainBench and validate its generalization capabilities on the external WCEB benchmark.
3.1 WebMainBench
@@ 3.1.1 Dataset Overview
WebMainBench comprises 7,887 carefully annotated web pages, each containing: (1) the raw HTML document, (2) ground-truth Main-HTML identified by human annotators as a valid DOM subtree, (3) the corresponding Markdown representation, and (4) rich metadata including language, content type, difficulty level, and structural characteristics. This comprehensive annotation enables fine-grained performance analysis across diverse web content types. An example data entry is shown in Appendix Figure 9.
@@ 3.1.3 WebMainBench-Structured: A Subset for Structured Element Extraction
To rigorously evaluate structured element preservation during main content extraction, we curate WebMainBench-Structured, a focused subset of 545 pages selected from WebMainBench that are rich in mathematical formulas, code blocks, and tables—elements critical for technical and educational content but often corrupted by traditional extraction methods. In addition to standard annotations, each page includes a groundtruth_content field containing ground-truth Markdown text with all structured elements faithfully preserved by human annotators, enabling precise element-level evaluation.
@@ 3.3.1 Main Content Extraction
Table 1: Mean ROUGE-N F1 on WebMainBench with different tracks
@@ 3.4 Generalization to WCEB
Table 3 shows that MinerU-HTML maintains strong performance on WCEB, achieving an overall score of 0.8002 and outperforming the strongest baseline, Trafilatura (0.7833). The strong performance across both WebMainBench and WCEB validates MinerU-HTML’s robustness and generalization capabilities across diverse web content.
@@ 5 Conclusion
We release MinerU-HTML, WebMainBench, and AICC to facilitate further research into semantic-aware extraction methods and their impact on pretraining data quality.
@@ 6.2 Document Formatting
Figure 9: An example data from WebMainBench. It includes the raw source, the ground-truth Main-HTML, its Markdown conversion, and a rich set of metadata for fine-grained analysis.
@@ 7.1 Benchmark Construction
Data Sampling. WebMainBench is constructed using a hybrid sampling strategy to ensure both broad representation and relevance. 90% of the samples are randomly drawn from the Common Crawl dataset to cover the long-tail web, while the remaining 10% are sampled from a list of top-ranking websites (Chinaz Alexa [^4] https://malexa.chinaz.com/) to include popular, professionally designed pages. The final benchmark is highly diverse, containing pages from 5,434 unique top-level and 5,904 unique second-level domains.